OpenStudy (anonymous):

The area of a triangle is 56.4 cm2 and the base is 12 cm. What is the height of the triangle?

OpenStudy (anonymous):

Need help.

OpenStudy (campbell_st):

the area formula is Area = 1/2 x base x height so you have 56.4 = 1/2 x 12 x height or 56.4 = 6 x height divide both sides of the equation by 6

OpenStudy (vianne):

Well you know the base which is 12 cm. Plus you know the area.. So all we need to do is divide it. 56.4 / 12 = 4.7 And that's your answer. :)

OpenStudy (anonymous):

Thank you!